### v1

# Extracting Probabilistic Knowledge from Large Language Models for Bayesian Network Parameterization

Abstract
Large Language Models (LLMs) have demonstrated potential as factual knowledge bases; however, their capability to generate probabilistic knowledge about real-world events remains understudied. This paper investigates using probabilistic knowledge inherent in LLMs to derive probability estimates for statements concerning events and their interrelationships captured via a Bayesian Network (BN). Using LLMs in this context allows for the parameterization of BNs, enabling probabilistic modeling within specific domains. Experiments on eighty publicly available Bayesian Networks, from healthcare to finance, demonstrate that querying LLMs about the conditional probabilities of events provides meaningful results when compared to baselines, including random and uniform distributions, as well as approaches based on next-token generation probabilities. We explore how these LLM-derived distributions can serve as expert priors to refine distributions extracted from minimal data, significantly reducing systematic biases. Overall, this work introduces a promising strategy for automatically constructing Bayesian Networks by combining probabilistic knowledge extracted from LLMs with small amounts of real-world data. Additionally, we evaluate several prompting strategies for eliciting probabilistic knowledge from LLMs and establish the first comprehensive baseline for assessing LLM performance in extracting probabilistic knowledge.
Extracting Probabilistic Knowledge from Large Language Models for Bayesian Network Parameterization
Aliakbar Nafar1, Kristen Brent Venable2,3, Zijun Cui1, Parisa Kordjamshidi1 1Michigan State University 2Florida Institute for Human and Machine Cognition (IHMC) 3University of West Florida {nafarali, cuizijun, kordjams}@msu.edu, bvenable@ihmc.org
1 Introduction
The potential of language models as sources for extracting factual knowledge has been demonstrated in several studies (Petroni et al., 2019; Roberts et al., 2020; AlKhamissi et al., 2022; Zhao et al., 2025). However, it remains unclear whether LLMs possess the ability to generate meaningful probabilistic estimates for events and their relationships based on their internal knowledge. We refer to an LLM’s capability to assign a specific probability to an uncertain proposition by utilizing its internal knowledge as probabilistic estimation. This contrasts with providing a confidence score for a concrete answer (Xiong et al., 2024) or solving a probabilistic problem that has a known numerical solution, a process referred to as probabilistic inference (Nafar et al., 2024b). For instance, consider the question, “What is the probability that a person who smokes cigarettes will develop cancer in their lifetime?” There is no definitive correct answer to infer directly from the question itself; however, a medical expert familiar with the literature might approximate 20%. Similarly, we expect an adept language model to produce a similar estimate.
Here, we evaluate the probabilistic estimation capabilities of LLMs such as GPT-4o (OpenAI et al., 2024), Gemini Pro 1.5 (Team et al., 2024), and Claude 3.5 (Anthropic, 2024), and utilize their internal knowledge to construct domain-specific Bayesian networks (BNs). BNs are a powerful method for representing uncertainty and dependencies between events. The reliability of inference in BNs hinges on the accuracy of conditional probability table (CPT) entries, which are typically obtained by collecting data or eliciting expert judgments. These approaches can be expensive or unattainable in domains where data is scarce (You et al., 2019; Longato et al., 2023) or expert time is limited (Das, 2008; Xiaoguang et al., 2019).
By extracting probabilistic knowledge of the LLMs, we explore whether they can serve the role of experts for parameterizing BNs. We assume the dependency structure within the BN is given, leaving the task of extracting such dependencies using LLMs for future research. Nonetheless, preliminary results indicate that LLMs show promise in accurately identifying network dependencies (Babakov et al., 2025). Given the structure of a BN, LLMs are required to predict a probability distribution for each node, conditioned on its parent nodes. In this work, we focus on discrete variables, so we query the LLM separately for each state of a node, prompting it to directly generate a numerical output representing the conditional probability given its parent nodes. After collecting these raw numerical outputs from the LLM, a normalization step is performed to ensure valid probability distributions.
In domains where only a small dataset is available, initial distributions estimated by LLMs can potentially function as expert-derived prior probabilities. We test this hypothesis by adjusting the LLM predictions with a small number of data samples, effectively applying a partial calibration to the model’s initial estimates. We demonstrate that this approach not only enhances the predictions of the LLMs but also surpasses purely data-driven methods and provides a superior prior compared to a uniform prior baseline, which is typically employed in the absence of additional information. Furthermore, our experiments indicate that even when the number of data samples is large, incorporating LLM priors improves the performance. These findings highlight the promise of leveraging LLMs as expert knowledge sources for probabilistic estimation across various real-world domains.
In summary, our contributions are as follows:
1) Comprehensive Evaluation of LLM Probabilistic Estimation with Real-World BNs. We provide the first large-scale empirical analyses of how LLMs quantify probabilities. We investigate differences in LLM accuracy across diverse BN domains and varying levels of conditional complexity, highlighting the strengths and limitations of LLMs as probabilistic knowledge bases.
2) LLM Predictions as Expert-Driven Priors. We demonstrate that the LLM predicted probabilities can serve as expert-derived priors and combining them with sampled data improves accuracy and sample efficiency compared to purely data-driven methods or using a uniform prior baseline.
[…]
2 Related Work
Parameter Learning. Within the field of Bayesian Networks, parameter learning has historically revolved around data-driven approaches (Ji et al., 2015) or expert elicitation (Mendes, 2014). Data-driven methods compute frequencies for each configuration of parent states, estimating conditional probabilities via frequentist or Bayesian methods. Expert elicitation is commonly used in domains where data is limited or unobserved for some variables (Woodberry et al., 2005; Constantinou et al., 2016). This work aims to see whether LLM-based probability queries can fill a similar role to experts.
Zero-shot Regression. Using LLMs for regression in a zero-shot setting is an emerging field, and work in this area is limited. Following Vacareanu et al. (2024), which shows that LLMs are capable regressors in a few-shot setting, Nafar et al. (2025) tests the regression capability of LLMs in a zero-shot setting (using internal knowledge) for realistic questions such as estimating the medical insurance cost based on age. However, they don’t use any probability estimation. Paruchuri et al. (2024) asks LLMs to calculate the probabilities for a range of values in a given distribution, but their dataset is limited to only 12 questions, and the queries are elementary with no conditional variables.
Confidence Elicitation. Confidence elicitation in LLMs has been studied in classification tasks, where a confidence score ranging from 0 to 1 is assigned to a discrete class label. Among these, Kadavath et al. (2022) treat the models as a white box and use their token probabilities to assess confidence. However, token likelihood merely indicates the model’s uncertainty about the next token (Kuhn et al., 2023), rather than the confidence in the text itself. Consequently, Xiong et al. (2024); Yang et al. (2024) treat the model as a black-box and use its generated confidence to solve classification datasets. However, confidence elicitation does not directly apply to scenarios requiring a probability distribution across multiple states. Moreover, while similar, confidence is inherently different from probability (Levine, 2024). Another stream of research, in which confidence aligns more closely with probability, prompts LLMs to produce binary probabilities for intermediate nodes, which are also created by LLMs, forming simplified graphical models tailored toward classification tasks (Huang et al., 2025; Feng et al., 2025). In contrast to these methods—dependent on smaller, synthetic graphs constructed explicitly for classification, our approach utilizes large-scale, real-world networks necessitating probability distributions across multiple states. Furthermore, while Feng et al. (2025) extracts hedge words from LLM outputs and subsequently trains a separate model to estimate corresponding probabilities, we directly prompt the LLM to provide explicit probability values.
Probabilistic Inference. Probabilistic inference is closely related to our task and can be considered a natural extension of probabilistic estimation. Saeed et al. (2021); Nafar et al. (2024a) fine-tune BERT-based language models to perform probabilistic inference, while Nafar et al. (2024b) utilizes prompt engineering techniques to enable LLMs to conduct probabilistic inference. However, in all these approaches, the explicit probabilities of sentences are either provided in the text or learned from the dataset during fine-tuning without any estimation derived from the prior internal knowledge.
3 Problem Definition
The main problem addressed in this paper is parameterizing a Bayesian Network given its structure. We formally define the problem as follows, given the structure of a Bayesian Network $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $V$ is the set of nodes (random variables) and $E$ is the set of edges (dependencies among variables), the goal is to estimate the parameters, that is, Conditional Probability Tables (CPTs) $\mathcal{G}_{\theta}$ of this network using LLMs. We investigate whether LLMs can be used to obtain these parameters. Specifically, the problem becomes one of posing textual queries describing a node’s parents and requesting a probability distribution for each row of the CPT. After this parameter assignment step, we compare the resulting distribution to ground-truth values of the original Bayesian Networks.
4 Methodology
4.1 Overview
We introduce two distinct frameworks for estimating the parameters of a BN. Our first approach, Extracting Probabilistic Knowledge (EPK), uses prompting techniques to derive probabilistic estimates. Instead of querying the entire probability distribution at once, we individually prompt the LLM for probabilities corresponding to each state of a node and subsequently normalize these values to ensure a valid distribution. Our second approach, Expert-Driven Priors (EDP), assumes the availability of a small empirical dataset and incorporates the LLM-generated predictions as expert-driven priors combined with the empirical data.
4.2 Extracting Probabilistic Knowledge (EPK)
Figure 1: BN Parameterization using EPK. The left side represents an input BN with its dependency structure. The right side shows how EPK uses LLMs to parameterize the BN. Given a node (e.g., Lung Cancer) and its parents, the LLM is prompted to estimate probabilities of its states. The node description is appended to each question to clarify the meanings and states. The raw probabilities generated by the LLM are then normalized to sum to one.
As depicted in Figure 1, the process starts with a prompt template that describes the node and its parents in a natural language format and instructs the LLM to provide a probability value. This step is necessary to establish the nodes’ meaning and possible values. The descriptions of these nodes, combined with the LLM instructions, are appended to each question presented to the LLM. Each question explicitly defines the states of the node of interest as well as the states of its parent nodes, and poses a probabilistic query based on these assigned states. Depending on the instruction, the LLM can produce the probability directly or articulate its reasoning before generating a probability value. In both settings, the final answer is extracted as a numerical probability from the output text.
In the questions posed to the LLM in Figure 1, providing descriptions for nodes and their states is optional when the node names and their corresponding values are straightforward. For example, the meanings of “Lung Cancer” and “Smoking” is intuitive, and the LLM can infer that their values are binary, based on the given assignments. However, in cases where the meaning or value sets of nodes are not immediately clear, explicit descriptions are essential. For instance, a node named “X1” must have a clearly stated meaning, such as: “Represents a lack of supervision and policy guidance, which may lead to the use of unqualified oil. This node can take on the values True or False.” Similarly, while the meaning of the “Construction Year” node is self-evident, its possible states are ambiguous. Providing a description like “This node indicates the time period in which the building was constructed, with possible values being 1930-1955, 1955-1960, 1960-1968, 1968-1975, and 1975-1980.”
Since the raw numeric outputs may not sum to one, they are normalized to form a valid distribution over the node’s states. For a node with $m$ states, the model might produce values $p_{1},p_{2},\dots,p_{m}$ that sum to $S=\sum_{i=1}^{m}p_{i}$. To convert these values into a valid probability distribution, we divide each one by $S$. This normalization step can be interpreted as taking the ratio of each state’s assigned likelihood relative to the sum of all states, effectively preserving the proportions while enforcing a valid distribution.
4.3 Expert-Driven Priors (EDP)
Expert opinion and large datasets are either cost-prohibitive or difficult to obtain in many practical scenarios. However, collecting a small number of real observations may still be feasible. When only limited data is available, incorporating expert prior knowledge can help offset the shortage of empirical data, thereby improving the resulting probability distributions. We propose that LLMs can fill the role of experts by providing these prior distributions based on their knowledge. In our approach, EDP, we combine the LLM-derived probabilities with the empirical distribution estimated from these few samples by utilizing two approaches: linear pooling (Gneiting and Ranjan, 2013) and using priors as pseudocounts (Zhai and Lafferty, 2001). For linear pooling, let $q_{1},\dots,q_{m}$ denote the probabilities assigned by the LLM (after normalization) for a particular node with $m$ possible states under a specific parent configuration, and let $r_{1},\dots,r_{m}$ represent the empirical probabilities derived using maximum likelihood estimation, obtained from the collected sample data. To form a final distribution, we apply a linear combination where each final probability is given by $p_{i}=\alpha q_{i}+(1-\alpha)r_{i}$ for $i=1,\dots,m$. Here, $\alpha\in[0,1]$ is a parameter that determines the weight of LLM-derived distribution relative to the empirical one. In the approach where we use the priors as pseudocounts, we translate the LLM-derived probabilities $q_{1},\dots,q_{m}$ into pseudocounts representing prior knowledge by introducing a confidence hyperparameter $C$ and computing pseudocounts $C\cdot q_{i}$ for each state $i$. Then, we combine these pseudocounts with observed counts $n_{i}$ from collected data, obtaining updated probabilities $p_{i}=(C\cdot q_{i}+n_{i})/(C+\sum_{j=1}^{m}n_{j})$ for $i=1,\dots,m$. The hyperparameters $\alpha$ and $C$, which determine the effectiveness of priors, are set to be the reciprocal of data sample numbers, reflecting the intuition that their effect should diminish with more data.
5 Experiments
5.1 Dataset of Eighty Bayesian Networks
Our experiments use bnRep (Leonelli, 2025), a publicly available collection of BNs spanning diverse domains such as medicine and engineering. These BNs vary in size, ranging from 5 to 50 nodes. Each BN in bnRep has an associated publication detailing its construction, data sources, and domain-specific context. Most of these papers were published between 2020 and 2024, making them well-suited for LLMs incorporating recent knowledge. Each BN is accompanied by ground-truth parameters, which enable us to evaluate our predictions. We preprocess bnRep to select only discrete BNs and eliminate BNs with missing entries in CPTs, yielding a final dataset of 80 networks. We extract node definitions and states from each BN’s repository paper. Additional details, including preprocessing steps, are provided in Appendix A.
5.2 Metrics, and Baselines
As of LLMs, we utilize GPT-4o and its mini variant (OpenAI et al., 2024) in most of our experiments, along with Claude Sonnet 3.5 (Anthropic, 2024) and Gemini-Pro 1.5 (Team et al., 2024). We evaluate the LLM’s estimated BN parameters using Kullback–Leibler (KL) divergence (Kullback and Leibler, 1951) compared to ground-truth parameters in bnRep. Specifically, we report the BN KL divergence, defined as the KL divergence computed over the BN variables’ joint distribution, evaluating the resulting BN’s overall quality. For details on the LLMs and their hyperparameters, see Appendix B, and for the explanation of KL divergence and BN KL divergence calculation, refer to Appendix C.
We evaluate our methods against multiple baselines. All outputs from these baselines are normalized as necessary to ensure valid probability distributions. These baselines are: (1) Random number generator; (2) Uniform baseline generating equal probabilities for each row of the CPT, providing basic, uninformed estimations; (2) LLM (Random) baseline involving intentionally incorrect queries, where the original variable names are randomly replaced. This is done to assess whether LLMs utilize the content of the provided questions to generate their answers. For instance, we query the LLM "What is the probability that construction time is true given that lung cancer is true?" instead of the correct question regarding smoking; (3) LLM (No Context) baseline in which queries are presented without contextual explanations, exploring scenarios where the node meanings and number of states cannot be directly inferred from the in-context information; (4) LLM (Token Probability) baseline, which directly uses the LLM’s internal probabilities assigned to tokens representing node states (e.g., probability of generating the token "True"), rather than explicitly generated numerical probabilities extracted from the model’s textual responses; (5) LLM Full Dist baseline wherein the LLM is prompted for the complete distribution of the node corresponding to a specific combination of parent states (a CPT row), and the response is received as a tuple of numbers. The order of the states is given in the prompt; and (6) MLE-# is a statistical baseline obtained by maximum likelihood estimation using data sampled from the ground-truth BN where the BN is sampled exactly # times using forward sampling. For results obtained using other sampling methods, refer to the Appendix D.
5.3 Can LLMs Estimate Probabilities Using their Internal Knowledge?
Figure 2: Boxplot showing the distribution of BN KL divergence values across 80 unique BNs for various models, sorted by their median KL divergence. Lower values indicate better alignment with ground truth CPTs.
To evaluate how EPK compares to other baseline models, we analyze the distributions of BN KL divergence across all 80 BNs, as depicted in Figure 8. As expected, the parameter estimation of statistical models (MLE-#) progressively improves as the size of the sample data increases. Models trained on minimal data demonstrate notably poor performance, similar to random guessing. Interestingly, the worst performing models, performing worse than random, are the “Token Probability” models. This aligns with previous research, which found that raw token probabilities from LLMs alone are insufficient for effective uncertainty/probability estimation (Xiong et al., 2024) and require additional processing steps like fine-tuning (Tao et al., 2024).
The next weakest results are observed among the random generators, including GPT-4o (Random) and GPT-4o-mini (Random), next to the baselines that do not observe the correct context for nodes and their states, GPT-4o (No Context) and GPT-4o-mini (No Context). We observe that LLM (Random) models slightly surpass the outputs of the Random number generator baseline. However, these improvements only reflect the non-uniformity of random number generation by LLMs, influenced by factors such as text-generation sampling methods and model architecture choices (Hopkins et al., 2023). Of all the baseline models, the uniform predictor performs best. This result aligns with information theory, which suggests that, in the absence of knowledge, a uniform distribution naturally provides the lowest KL divergence based on uncertainty (Cover and Thomas, 2006).
Lastly, both the LLM (Full Dist) baseline and EPK outperform the uniform baseline. However, while the LLM (Full Dist) baseline yields informed estimates, it consistently falls short of EPK with a higher median KL divergence and a greater standard deviation across the BNs. The MLE-# model requires at least 30 samples to match the median KL divergence of EPK. Even then, its standard deviation remains higher, indicating that EPK is more robust. Overall, these results demonstrate the capability of LLMs to provide meaningful probability distributions based on their internal knowledge.
5.4 Can LLMs’ probability estimations serve as Expert Priors?
Figure 3: Distribution of BN KL divergence across 80 BNs using GPT-4o (EDP) and Uniform predictions as priors combined with empirical data via the linear pooling method. The data sample size is shown in parentheses. The MLE-# models only use data samples, and the EPK models only use LLMs for probability estimations.
We evaluate the effectiveness of EDP, combining LLM-derived distributions with empirical data, using two methods: priors as pseudocounts and linear pooling. Figure  3 displays the distributions of BN KL divergences obtained by combining GPT-4o and Uniform priors with various sample sizes of data with the linear pooling method. We chose GPT-4o because it has the best performance among the evaluated LLMs. At the same time, the Uniform model serves as the best-performing baseline in our experiments and is typically used in the absence of prior information. The method using priors as pseudocounts shows the same trend, and its results are presented in Appendix D to save space.
Regardless of the combination method, GPT-4o predictions consistently outperform the Uniform model baseline, proving its use as a better prior. The advantage of incorporating GPT-4o priors is most notable at smaller sample sizes, i.e., 3 to 10 samples. The combination of even minimal data in EDP significantly outperforms EPK and the MLE model with 30 data samples. Additionally, EDP improves the median KL divergence even when abundant data, 1000 or more samples, is used. It also effectively reduces the standard deviation of the predictions, enhancing model robustness. This improvement in median KL divergence at large sample sizes occurs because nodes with unlikely parent combinations rarely receive data. This is supported by the absence of KL divergence improvement at high sample sizes with alternative methods that separately sample data for each CPT row (See Appendix D). However, these alternative, unrealistic methods still yield decreased standard deviation and improved robustness. These findings confirm that incorporating LLM predictions as expert-driven priors is beneficial and enhances performance and robustness.
6 Discussion
6.1 Quality of Individual Distributions Versus the Entire Bayesian Network
In our experiments, we employed the BN KL Divergence metric to evaluate the quality of the predicted BNs. However, this metric assesses the entire BN, meaning that a few poorly predicted nodes might disproportionately influence the evaluation. To address this limitation, we also analyzed the CPT KL Divergence, which computes the average KL divergence across all individual CPT rows within each BN. This alternative measure evaluates the quality of individual distributions rather than the BN as a whole. Using CPT KL Divergence, we observed that the overall trends of our results remained consistent. Additional diagrams illustrating these findings are provided in Appendix D.
6.2 Performance Variations Among LLMs Across Different Bayesian Networks
Among the evaluated LLMs, GPT-4o consistently exceeds the performance of other LLMs, though specific models perform better within specialized domains. For instance, Claude 3.5, Gemini-pro, and GPT-4o achieve the best results on BNs related to engineering, business, and medical domains, respectively. Furthermore, there are inherent differences in prediction behavior among these LLMs, likely attributed to their respective training methodologies. Specifically, Claude 3.5 performed best on BNs with low entropy probabilities, but showed the poorest performance on BNs with high entropy probabilities among the LLMs, indicating an overly confident prediction behavior. In contrast, Gemini-pro showed the opposite trend, whereas GPT-4o had a more balanced prediction profile.
6.3 Prompt Styles and Chain-of-Thought
In our experiments, consistent with previous research (Xiong et al., 2024), we observed notable volatility in token probabilities, with significant variations depending on preceding tokens. However, their behavior remained robust when prompting the LLMs to verbally generate probabilities as textual output. Experimental results remained consistent whether we instructed the model to explain its reasoning before providing probabilities or to output probabilities directly. These findings align with prior research on zero-shot regression (Nafar et al., 2025), indicating that chain-of-thought reasoning did not affect results. Furthermore, other prompting styles or additional contextual information taken from relevant Bayesian Network publications also did not influence the results. Our experiments suggest that although LLMs clearly possess and can articulate probabilistic knowledge, this capability remains consistent irrespective of explicit explanatory reasoning.
6.4 Handling Larger Parent Sets and States
Intuitively, it makes sense that LLMs may struggle to provide informed predictions for more complex queries involving nodes with many parent nodes or states. We use the CPT KL Divergence metric to assess predictions among these nodes, which averages the KL divergence across all individual CPT rows rather than evaluating the entire BN. In our experiments with realistic BNs, LLM performance in these scenarios still surpassed our baseline models. LLMs consistently outperformed baselines in queries involving up to 5 parent nodes. Additionally, the LLMs performed better than baselines for nodes with 2 or 3 states. Nodes exceeding 3 states are rare in realistic BNs. Only 4 BNs had nodes restricted to 4 states, whereas 11 featured nodes with 5 states. A notable exception was the ’DustExplosion’ BN, which contained nodes with both 4 and 5 states. This network, designed to predict explosion probabilities in industrial environments, proved challenging for the LLMs. Apart from this isolated case, the LLMs consistently outperformed baseline methods. For detailed results demonstrating the performance with varying numbers of parent nodes and states, refer to Appendix E.
6.5 Trade-off Between EPK and Full Distribution Queries
Our experiments show that EPK consistently outperforms directly requesting the complete distribution (Full Dist), achieving lower median KL divergence and improved stability. Although the Full Dist method is more straightforward and requires fewer queries, it leads to less precise and more variable estimates. Additionally, our proposed approach remains cost-effective by avoiding extensive reasoning or explanations, as most LLM-related costs derive from text generation. Thus, we recommend EPK for applications prioritizing accuracy and consistency, while Full Dist might be suitable in scenarios where cost saving is paramount.
6.6 Implications of the EDP method for Automated BN Construction
The demonstrated capability of LLMs, particularly GPT-4o, in the EDP method has significant implications for automating BN construction. Traditionally, parameterization of Bayesian networks relies heavily on expert input, making the process labor-intensive and costly. However, cost is not the only limitation associated with expert-driven approaches. In some cases, experts may not even be available, and when they are, their expertise must be evaluated (Hald et al., 2016) or aggregated with the opinion of other experts (McAndrew et al., 2021) before their views can be utilized. Also, expert opinions are inherently susceptible to errors such as overconfidence and bias (Soll et al., 2024). Utilizing an LLM proficient across diverse domains holds enormous potential for advancing automation in BN parameterization.
6.7 Combining LLMs’ Priors with Small Data for Extraction of Probabilistic Knowledge
In Section 5.4, we showed the potential of using LLM predictions in place of expert-driven priors for constructing BNs. Incorporating even a small amount of data, for example, just 3 samples, together with LLM predictions, yields a lower BN KL divergence than models trained purely on 30 data samples or those relying solely on LLM predictions. These results highlight a promising application wherein LLMs use minimal external data points to rapidly refine their predictions for probabilistic queries. Such a small amount of data could be supplied to the LLM in various ways, such as being obtained online by querying information from publicly available sources (e.g., occurrences of lung cancer among smokers). Combining LLMs’ priors with data in the EDP approach not only improves data efficiency and outperforms uniform baseline methods but also provides an exciting possibility for real-time improvement of LLM-generated probabilistic queries with minimal data input.
7 Conclusion and Future Work
In conclusion, this study introduces a novel pipeline for parameterizing BNs using LLMs and establishes the first comprehensive framework for evaluating the probabilistic knowledge of LLMs with real-world probabilistic BNs. We systematically assess various BNs and demonstrate that modern LLMs can reliably produce conditional probabilities, outperforming baselines and purely data-driven approaches, especially in a low-data regime. Importantly, our evaluation framework highlights the probabilistic estimation capabilities of LLMs, offering insights into their internal knowledge, demonstrating their value as expert-based priors, and facilitating automated BN construction.
Future work will focus on advancing toward a fully automated framework for Bayesian Network construction. In this regard, the key challenge lies in automating the structure learning component. Although preliminary efforts have been made in this area, there is significant potential to create an end-to-end pipeline using LLMs that generates a BN structure, parameterizes it, and systematically evaluates its performance.
Limitations
Our analysis demonstrates clear advantages in employing LLMs for Bayesian network parameter estimation. While KL divergence is effective for comparative evaluation within our context, it primarily indicates relative improvements rather than quantifying specific practical benefits for individual downstream tasks, which often require tailored metrics, such as assessing the accuracy of query probabilities within a defined precision (e.g., a 5% threshold). Incorporating domain expert judgments to evaluate the BNs and design specific metrics for each BN would benefit our work, but the associated costs and resource requirements substantially limit this approach. Expert evaluations would realistically be feasible only for a select subset of Bayesian networks rather than all 80 networks analyzed, limiting the comprehensiveness and comparative scope of the evaluation currently achievable through automated methods.
Additionally, our study utilizes only state-of-the-art models because smaller, more economical models performed poorly in the task of probabilistic knowledge estimation. Although our approach avoids fine-tuning and thus has lower costs than fine-tuning-based methods, the inference costs remain relatively high among inference-only approaches, potentially posing challenges in particularly cost-sensitive scenarios.
Appendix A Dataset Pre-processing
A.1 bnRep Dataset Overview
The bnRep dataset (Leonelli, 2025) is an open-source collection designed to facilitate research, teaching, and practical applications related to Bayesian Networks (BNs), addressing the significant shortage of comprehensive BN repositories. Implemented as an R package, bnRep includes over 200 well-documented Bayesian Networks sourced from more than 150 academic publications, mainly recent studies published from 2020 onwards. Each BN entry has accompanying characteristics extracted from the original publications. These characteristics are described below:
  • Name: A short identifier for the Bayesian network.
  • Type: The network’s type of random variables (discrete, continuous, mixture).
  • Structure: Indicates how the network’s structure was obtained.
    • Knowledge: The structure is built from well-established domain knowledge.
    • Data: The structure was learned from a dataset.
    • Fixed: A predefined structure that is neither purely elicited from experts nor learned from data (often a canonical or standard network).
    • Synthetic: The structure was generated artificially (e.g., for algorithm testing).
    • Expert: The structure is directly elicited from domain experts.
    • Mixed: The structure is derived through a combination of sources.
  • Probabilities: Indicates how the CPTs were obtained:
    • Data: Parameters estimated from empirical data.
    • Knowledge: Parameters derived from well-established theoretical or domain-specific information.
    • Mixed: A combination of data-based estimation and expert input.
    • Synthetic: Artificially generated parameters for testing or demonstration.
    • Expert: Parameters directly elicited from domain experts.
  • Graph: Describes any special structural characteristic of the network graph. For example:
    • Generic: No particular restriction or canonical form.
    • Naive Bayes: A star-shaped structure often used for classification tasks.
    • Reverse Naive Bayes: The class label is modeled as a child of all other variables, reversing the direction of edges in the standard Naive Bayes structure.
    • K-Dep: Each feature depends on the class and up to K other features.
    • Tree: A graph with each node having exactly one parent (except the root).
    • Reverse Tree: A tree with reversed edges, placing the class node at the leaves.
    • TAN: an extension of Naive Bayes that allows each variable to have one additional parent, forming a tree among the predictors for greater flexibility.
  • Area: The domain of the BN (e.g., Medicine, Engineering, Environmental Science).
  • Nodes: The total number of random variables (nodes) in the Bayesian network.
  • Arcs: The total number of directed edges (arcs) in the network.
  • Parameters: The total number of probability entries in the CPTs.
  • Avg. Parents: The average number of parent nodes per variable.
  • Max Parents: The maximum number of parents any single node has in the network.
  • Avg. Levels: The average number of discrete states (levels) per node.
  • Max Levels: The maximum number of states among all nodes in the network.
  • Average Markov Blanket: The average size of the Markov blanket for each node, which consists of the node’s parents, children, and the children’s other parents.
  • Year: The year of publication associated with the BN’s reference.
  • Journal: The venue where the Bayesian network was published or described.
  • Reference: The bibliographic reference describing the BN in detail.
A.2 BN Selection and Filtering
To utilize the bnRep dataset, we first converted the Bayesian networks from the R package into Python-compatible format using the pgmpy library (Ankan and Panda, 2015). Next, we filtered the networks, selecting only those containing discrete CPT values, as these comprise the majority (90%) of the BNs. Then, to improve practicality and computational efficiency, we excluded networks with more than 50 nodes, noting that only a few exceed this threshold. The most substantial reduction, however, came from removing BNs with incomplete CPT information. After applying all these criteria, we arrived at a final dataset of 80 BNs, which remains sufficiently large for our evaluation purposes.
A.3 Nodes’ Descriptions and States
As outlined in Section 4, querying the LLMs for each node state requires clearly defined descriptions for the nodes and their respective states. To acquire this information for each Bayesian Network (BN), we first retrieved the PDFs of the referenced documents that detail each BN. These references were initially provided by the bnRep dataset.
A Python script was developed to automate the download of relevant PDFs directly from Google Scholar using bibliographic information. Despite automation, several documents could not be downloaded automatically due to access restrictions or other limitations. These remaining documents were obtained manually, with some requiring specific permissions. Subsequently, we performed a manual verification to ensure the downloaded documents corresponded to their respective references.
We then implemented another Python script leveraging GPT-4o, which, given a PDF and the extracted nodes and states of the Bayesian Network (obtained through the pgmpy library), automatically generated a Python dictionary describing each node and its associated states. Although the majority of generated dictionaries were accurate and required minimal modifications, a few necessitated manual adjustments to ensure correct formatting and accuracy. An illustrative example of the prompt and the generated dictionary for the Bayesian Network “algalactivity2” is provided in Figure 4.
Figure 4: Example of node descriptions and state extractions generated by GPT-4o for the Bayesian Network “algalactivity2”. The Python dictionary describes each node and the possible values it can assume based on context from the provided document.
Appendix B LLM’s Hyper-Parameters and Experimental Setup
In our experiments, we primarily evaluated three state-of-the-art Large Language Models: GPT-4o and its mini variant (OpenAI et al., 2024) versions “gpt-4o-2024-11-20” and “gpt-4o-mini-2024-07-18”, Claude Sonnet 3.5 (Anthropic, 2024) version “Claude 3.5 Sonnet 2024-10-22”, and Gemini-Pro 1.5 (Team et al., 2024) version “gemini-1.5-pro-002”. All models were interfaced using the LangChain framework (Chase, 2022), ensuring consistent integration and interaction.
A fixed generation “temperature” of $0.1$ was utilized to maintain minimal stochasticity in outputs. We initially explored the impact of sampling by performing up to five repeated samples per inference. However, we found that multiple samples did not meaningfully affect aggregate outcomes across the evaluated set of 80 Bayesian Networks, likely due to the low temperature setting. Therefore, subsequent experiments were conducted using a single sample per inference.
Token output lengths were not constrained, allowing the models to elaborate their reasoning freely. In instances where models produced responses that deviated from the required format or where output text generation was interrupted midway, additional prompts were provided until valid responses were obtained. The prompt structure used for eliciting responses is illustrated at the top of Figure 5. To get the token probabilities from our LLMs, we use the middle prompt template in Figure 5. OpenAI’s Chat API allows for extracting the top 20 most probable tokens. To get the full distribution of a node’s states at once using the Full Dist method, we use the prompt template at the bottom of Figure 5.
Figure 5: Prompt templates used for eliciting probabilistic responses from LLMs. The top panel shows the main prompt template that asks the LLM to provide a probabilistic answer along with explicit reasoning. The middle panel shows the modified prompt template designed specifically to extract token-level probabilities by instructing the LLM to generate only the most probable state without additional context or explanation. The bottom panel shows the prompt template used for the Full Dist method, which requests the model to output the entire probability distribution over a node’s possible states in a single response.
Appendix C KL Divergence
C.1 Kullback Leibler (KL) Divergence Overview
Kullback Leibler (KL) divergence measures how one probability distribution $p$ diverges from a reference distribution $q$. For a discrete random variable $X$,
$D_{\mathrm{KL}}\!\bigl{(}p(X)\,\|\,q(X)\bigr{)}=\sum_{x\in\mathcal{X}}p(x)\,\log\frac{p(x)}{q(x)}.$
For continuous variables the sum is replaced by an integral. KL divergence satisfies $D_{\mathrm{KL}}(p\|q)\geq 0$ and equals zero iff $p=q$. It is asymmetric ($D_{\mathrm{KL}}(p\|q)\neq D_{\mathrm{KL}}(q\|p)$ in general). Intuitively, it quantifies the expected extra amount of information (in nats or bits) required to encode samples from $p$ using a code that is optimal for $q$.
C.2 BN KL Divergence Calculation
Calculating the KL divergence over the entire Bayesian network can be computationally expensive. To address this, we decompose it into a weighted sum of CPTs, where the weights correspond to the probabilities of the parent nodes, as illustrated in Figure 6. This approach simplifies computations in practice, since we only need to infer the probabilities of parent nodes using variable elimination.
BN KL Divergence Decomposition into a Sum of Local KL Divergences Let $p(\mathbf{x})$ and $q(\mathbf{x})$ be two BNs over the same variables $\{X_{1},\dots,X_{n}\}$ with common structure. Each factorizes as $p(\mathbf{x})=\prod_{i=1}^{n}p\ \bigl{(}x_{i}\mid\mathrm{Pa}_{i}\bigr{)},\quad q(\mathbf{x})=\prod_{i=1}^{n}q\ \bigl{(}x_{i}\mid\mathrm{Pa}_{i}\bigr{)},$ where $\mathrm{Pa}_{i}$ are the parents of $X_{i}$. We want to show: $D_{\mathrm{KL}}(p\,\|\,q)=\sum_{i=1}^{n}\sum_{\mathrm{pa}_{i}}p(\mathrm{pa}_{i})\,D_{\mathrm{KL}}\!\bigl{(}p(X_{i}\mid\mathrm{pa}_{i})\,\|\,q(X_{i}\mid\mathrm{pa}_{i})\bigr{)}.$
Derivation.
$D_{\mathrm{KL}}(p\,\|\,q)$ $=\sum_{\mathbf{x}}p(\mathbf{x})\,\log\frac{p(\mathbf{x})}{q(\mathbf{x})}=\sum_{\mathbf{x}}p(\mathbf{x})\,\log\frac{\displaystyle\prod_{i=1}^{n}p(x_{i}\mid\mathrm{Pa}_{i})}{\displaystyle\prod_{i=1}^{n}q(x_{i}\mid\mathrm{Pa}_{i})}=\sum_{\mathbf{x}}p(\mathbf{x})\sum_{i=1}^{n}\log\frac{p(x_{i}\mid\mathrm{Pa}_{i})}{q(x_{i}\mid\mathrm{Pa}_{i})}$ $=\sum_{i=1}^{n}\sum_{\mathbf{x}}p(\mathbf{x})\,\log\frac{p(x_{i}\mid\mathrm{Pa}_{i})}{q(x_{i}\mid\mathrm{Pa}_{i})}=\sum_{i=1}^{n}\sum_{\mathrm{pa}_{i}}p(\mathrm{pa}_{i})\sum_{x_{i}}p(x_{i}\mid\mathrm{pa}_{i})\,\log\frac{p(x_{i}\mid\mathrm{pa}_{i})}{q(x_{i}\mid\mathrm{pa}_{i})}$ $=\sum_{i=1}^{n}\sum_{\mathrm{pa}_{i}}p(\mathrm{pa}_{i})\,D_{\mathrm{KL}}\!\bigl{(}p(X_{i}\mid\mathrm{pa}_{i})\,\|\,q(X_{i}\mid\mathrm{pa}_{i})\bigr{)}.$
Example.
Let $A,B,C\in\{0,1\}$ within the network $A\!\to\!B\!\to\!C$. We compute $D_{\mathrm{KL}}(p\|q)$ as follows: 1. Substitute the factorizations: $D_{\mathrm{KL}}(p\|q)=\sum_{a,b,c}p(a,b,c)\,\log\frac{p(a,b,c)}{q(a,b,c)}.=\sum_{a,b,c}p(a,b,c)\,\log\frac{p(a)\,p(b\mid a)\,p(c\mid b)}{q(a)\,q(b\mid a)\,q(c\mid b)}.$ 2. Separate logs: $=\sum_{a,b,c}p(a,b,c)\Bigl{[}\log\frac{p(a)}{q(a)}+\log\frac{p(b\mid a)}{q(b\mid a)}+\log\frac{p(c\mid b)}{q(c\mid b)}\Bigr{]}.$ 3. Split the sum: $=\underbrace{\sum_{a,b,c}p(a,b,c)\,\log\frac{p(a)}{q(a)}}_{A}+\underbrace{\sum_{a,b,c}p(a,b,c)\,\log\frac{p(b\mid a)}{q(b\mid a)}}_{B}+\underbrace{\sum_{a,b,c}p(a,b,c)\,\log\frac{p(c\mid b)}{q(c\mid b)}}_{C}.$ 4. Marginalize: $A=\sum_{a}p(a)\,\log\frac{p(a)}{q(a)},\quad B=\sum_{a}p(a)\sum_{b}p(b\mid a)\,\log\frac{p(b\mid a)}{q(b\mid a)},\quad C=\sum_{b}p(b)\sum_{c}p(c\mid b)\,\log\frac{p(c\mid b)}{q(c\mid b)}.$ 5. Recognize KL pieces and combine: $D_{\mathrm{KL}}(p\|q)=D_{\mathrm{KL}}\!\bigl{(}p(A)\|q(A)\bigr{)}+\sum_{a}p(a)\,D_{\mathrm{KL}}\!\bigl{(}p(B\mid a)\|q(B\mid a)\bigr{)}+\sum_{b}p(b)\,D_{\mathrm{KL}}\!\bigl{(}p(C\mid b)\|q(C\mid b)\bigr{)}.$
Figure 6: KL divergence decomposition into local components for Bayesian networks.
Appendix D Additional Diagrams
The results of the priors as pseudocounts method are shown in Figure 7. The trend and conclusions drawn from this figure are the same as the linear pooling method shown in Figure 3. Figure 8 shows the results of Figure 2 but with CPT KL divergence metric. The results of EDP with data sampled for each row of the CPT by # times the number of states of the node are shown in Figures 9 and 10.
Figure 7: Distribution of BN KL Divergence across 80 Bayesian Networks using GPT-4o (EDP) and Uniform predictions as priors combined with empirical data via the priors as pseudocounts method.
Figure 8: Boxplot showing the distribution of BN KL Divergence values across 80 unique Bayesian Networks for various models and baselines. Models are sorted by their median KL Divergence, with lower values indicating better alignment with ground truth CPTs.
Figure 9: Distribution of BN KL Divergence across 80 Bayesian Networks using GPT-4o (EDP) and Uniform predictions as priors combined with empirical data via the priors as pseudocounts method.
Figure 10: Distribution of BN KL Divergence across 80 Bayesian Networks using GPT-4o (EDP) and Uniform predictions as priors combined with empirical data via the linear pooling method.
Appendix E Differnt number parnets and states
Figures 11, 12, 13, 14 and 15 show the CPT KL divergence for nodes with 2, 3, 4, 5 and 6 states respectively. Figures 16, 17, 18, 19, 20, 21, 22 and 23 show the CPT KL divergence for nodes with 0, 1, 2, 3, 4, 5, 6 and 7 parents respectively.
Figure 11: CPT KL divergence for nodes with 2 states.
Figure 12: CPT KL divergence for nodes with 3 states.
Figure 13: CPT KL divergence for nodes with 4 states.
Figure 14: CPT KL divergence for nodes with 5 states.
Figure 15: CPT KL divergence for nodes with 6 states.
Figure 16: CPT KL divergence for nodes with 0 parents.
Figure 17: CPT KL divergence for nodes with 1 parents.
Figure 18: CPT KL divergence for nodes with 2 parents.
Figure 19: CPT KL divergence for nodes with 3 parents.
Figure 20: CPT KL divergence for nodes with 4 parents.
Figure 21: CPT KL divergence for nodes with 5 parents.
Figure 22: CPT KL divergence for nodes with 6 parents.
Figure 23: CPT KL divergence for nodes with 7 parents.